OpenStudy (anonymous):

Consider the function f(x) = x - x^2 on the interval (0,1). Which of the following statements is true? a. f is increasing b. f is decreasing c. f is constant d. f is neither increasing, nor decreasing, nor constant

OpenStudy (anonymous):

at the risk of repeating myself, if you find the vertex you will see that the \(x\) coordinate is \(\frac{1}{2}\) which is smack dab in the middle of your interval

OpenStudy (anonymous):

I think the answers to choose from are a bit dubious...

OpenStudy (anonymous):

So is it constant?

OpenStudy (anonymous):

no it is a quadratic polynomial, so it is never constant

OpenStudy (anonymous):

So if it is in the middle of the interval it would be D?

OpenStudy (anonymous):

yes, increases on \((-\infty, \frac{1}{2})\) and decreases on \((\frac{1}{2}, \infty)\)

OpenStudy (anonymous):

Ok! I get it now. Thank you :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

So D is also not really true right? :) They probably meant "None of the above statements is true"

OpenStudy (anonymous):

Yeah whatever gets me the credit lol